Use technology to find the​ P-value for a​ two-tailed test with nequals11 and test statistic t equals 3.075 . ​P-valuealmost equals nothing ​(Round to four decimal places as​ needed.)
Answer #1

solution:-

given that

test statistic t = 3.075

and n = 11

degree of freedom df = n - 1 = 11 - 1 = 10

the P-value for a two tailed test is

=> P-value = 0.0117